OpenStudy (anonymous):

sqrt 5(8+3 sqrt 6) 8 sqrt5 + 3 sqrt30 5 sqrt 10 130 None of the above

OpenStudy (ash2326):

Is this your question ? \[5\sqrt{8+3\sqrt 6}\]

OpenStudy (anonymous):

no its √5 (8 + 3√6)

OpenStudy (anonymous):

@ash2326

OpenStudy (ash2326):

OK \[\sqrt 5 (8+3\sqrt 6)\] Is this correct?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

Open the bracket like this \[A(B+C)=A\times B+A \times C\] \[\sqrt 5 (8+3\sqrt 6)=\sqrt 5\times 8 +\sqrt 5\times 3 \sqrt 6\] =\[8\sqrt 5+3(\sqrt 5\times \sqrt 6)\] Can you simplify this ? Remember \[\sqrt A \times \sqrt B= \sqrt{AB}\]

OpenStudy (ash2326):

@Surferbumm Do you get this?